OpenStudy (anonymous):

A grocer wants to make a 10-pound mixture of cashews and peanuts that he can sell for $3.29 per pound. If cashews cost $5.60 per pound and peanuts cost $2.30 per pound, how many pounds of each must he mix?

OpenStudy (anonymous):

Let c and p be the mixture weight of the cashews and peanuts respectively. Solve the following simultaneous equations for c and p.\[\left\{c+p=10, \frac{560}{100}c+\frac{230}{100}p=\frac{329}{100}*10\right\} \]

OpenStudy (anonymous):

\[\{c=3 \text{ lbs},p=7\text{ lbs}\} \]